dy/dx =x(ex2 +2)/6y2

1
Expert's answer
2021-03-25T14:06:04-0400

"{dy \over dx} ={x(e^{x^2} +2) \over 6y^2}"

"\int6y^2dy=\int xe^{x^2} dx+ \int 2x dx"


To integrate "\int xe^{x^2} dx"


Let "{x^2} =u, then"


"{du \over dx}=2x"


"{du \over 2} = xdx"


"{1 \over 2} \int e^ u du= {e^{x^2} \over 2} +c"


"\therefore \int6y^2dy=\int xe^{x^2} dx+ \int 2x dx"


"2y^3={e^{x^2} \over 2} + x^2+c"


"y=" "({e^{x^2} \over 4} + {x^2 \over 2} + {c \over 2})^{1 \over 3}"